# Exploring Holomorphic Retracts

### v6

@@ 1. Introduction
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose that $\partial D$ is convex near $p\in\partial D$ but with the boundary not being $\mathbb{C}$-extremal at $p$. Let $Z$ be an arbitrary one-dimensional retract of $D$ through the origin along the direction of $p$. Then $Z$ can be realized as the graph of a holomorphic map on the linear retract ${\rm span}(p)\cap D$, taking values in a bounded balanced pseudoconvex domain within the face $F_{\mathbb{C}}(p,\partial D)$ where $F_{\mathbb{C}}(p,\partial D)$ is the complex face of $\overline{D}$ at the boundary point $p$.
Next, we wish to go beyond the purview of Vesentini’s result. To do this by starting with the simplest step, we retain the assumption of balancedness; also, as the most natural class of domains in several complex variables are pseudoconvex domains, we ask if convexity in the above theorem of Vesentini can be relaxed to pseudoconvexity with no further assumptions about convexity as in the above theorem and extremality conditions dropped as well. Towards this, we have the following result wherein we assume boundedness. It must be noted that dropping boundedness leads to the failure of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ a retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$.
Relegating some relevant observations here to section 3, we record here a corollary for an important class of domains. Indeed, when $D$ as in the above theorem (without any extremality condition about any boundary point), is homogeneous i.e., ${\rm Aut}(D)$ acts transitively on $D$, then of-course we may glean a characterization of all retracts of such a domain i.e., retracts not passing through the origin as well.
@@
Owing to the assumption of being balanced, we recall that the domains as in this corollary are the much studied bounded symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV alongwith two exceptional cases, all of whose automorphism groups are well-known (Hua [Hua]). This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.7 for the $l^{p}$-ball for $p<\infty$ and theorem 1.8 for the $p=\infty$ case much earlier above; and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult. Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$. As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection. As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem, wherein which we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.13. It must first be noted that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs. This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc. Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains and with a simpler proof following the recent work [BBMV-union]. As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll], where it was shown that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like. We shall address this here, to the extent we can.
@@
Let $B_{1},B_{2}$ be a pair of bounded balanced strictly convex domains possibly in different complex Euclidean spaces. Then, every non-trivial retract of $D=B_{1}\times B_{2}$ through the origin is a holomorphic graph over such a linear subspace which itself is a retract of $D$. Infact, every non-linear retract of $D$ is the graph in $D$, either of a $B_{2}$-valued holomorphic map over a (complex) linear subspace of $B_{1}$ or, of a $B_{1}$-valued holomorphic map over a linear subspace of $B_{2}$ which itself is a linear retract of the respective ball.
More specifically and precisely, let $B_{1}\subset\mathbb{C}^{N_{1}}$ and $B_{2}\subset\mathbb{C}^{N_{2}}$ be as above i.e., bounded balanced strictly convex domains. Let $D=B_{1}\times B_{2}$ and suppose without loss of generality that $\min\{N_{1},N_{2}\}=N_{1}$. Let $\mu_{1},\mu_{2}$ denote the Minkowski functionals of $B_{1},B_{2}$ respectively, whose convexity ensures that $\mu_{1},\mu_{2}$ are norms on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively. Let $Z$ be any non-trivial retract of $D$, $L=T_{0}Z$, the tangent space to $Z$ at the origin, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$. Then:
@@
    if $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$, then $Z$ is a complex linear subspace of $D$ which can be realized as the graph of a complex linear map over the linear subspace $\pi_{1}(D_{L})$ of $B_{1}$ (the ball whose dimension is $\min\{N_{1},N_{2}\}$):
@@
    is an isometry from the Banach space ($\pi_{1}(L)$, $\mu_{1}$) into the Banach space ($\mathbb{C}^{N_{2}}$, $\mu_{2}$) and therefore maps $B_{1}^{L}:=\pi_{1}(D_{L})$ which is the unit ball in $\pi_{1}(L)$ with respect to $\mu_{1}$, isometrically into $B_{2}$; $\pi_{1}(D_{L})$ itself is a linear retract of $B_{1}$,
@@
    if $\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$, then $Z$ is the graph of a $B_{2}$-valued holomorphic map of a linear subspace $(B_{1})_{L_{1}}$ of $B_{1}$, with $(B_{1})_{L_{1}}$ being a linear retract of $B_{1}$,
@@
    if $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$, then $Z$ is the graph of a $B_{1}$-valued holomorphic map of a linear subspace $(B_{2})_{L_{2}}$ of $B_{2}$, with $(B_{2})_{L_{2}}$ being a linear retract of $B_{2}$,
@@
We cannot trivially extend this theorem to the case of a product $D=B_{1}\times B_{2}\times\ldots\times B_{n}$ – where $B_{j}$’s are all bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition – just by replacing one of the factors in the above theorem by a product of $(n-1)$ of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{n-1})\times B_{n}$, appealing to the above theorem and using an inductive argument, owing to the fact that $B_{1}\times B_{2}\times\ldots\times B_{n-1}$ fails to satisfy the strict convexity condition stipulated in the above theorem as soon as $n>2$. However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$. As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3. Here we record an application to the union problem for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornaess – Sibony in their article [Fost]. Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
@@
Take $\Omega$ in the union problem to be any homogeneous bounded balanced domain. If $M$ is non-hyperbolic and the corank of its (infinitesimal) Kobayashi metric is one, then $M$ is biholomorphic to $\Omega_{L}\times\mathbb{C}$ where $\Omega_{L}$ is a linear retract of $\Omega$.
Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is all the more significant. Towards this, we report the following result.
@@
Let $D$ be a bounded balanced strictly convex domain in $\mathbb{C}^{N}$. Let $Z$ be a retract of $D$ (not necessarily passing through origin) and suppose at-least one of the retraction maps for it say $\rho$, is a submersion at the origin. Then $Z=\rho(D)$ is the graph of a holomorphic function over some linear subspace $D_{L}$ of $D$. Indeed, $D_{L}=L\cap D$ where $L=T_{p}Z$ and $p=\rho(0)$.
Now, we have for the most part focussed attention on balanced domains which being contractible may be regarded as being topologically trivial. To turn next to topologically non-trivial cases, one simple case to begin with in this direction is that of analytic complements of domains with a Stein neighbourhood basis; specifically, see lemma 8.1. Next, to consider the simplest domain lacking a Stein neighbourhood basis then, we are led to considering the standard domain in this context namely the Hartogs triangle given by
@@
We also contend that this is perhaps the simplest topologically non-trivial case to consider because it is biholomorphic to the product $\Delta\times\left(\Delta\setminus\{0\}\right)$; indeed the map $\phi(z,w)=(w/z,z)$, effects the aforementioned biholomorphic correspondence. Thus it suffices to determine the retracts of $\Delta\times\Delta^{*}$, which we do in the following theorem which leads to a complete characterization the retracts of the Hartogs triangle as well as its ‘analytic complements’.
@@
While retracts of $\Omega_{H}$ as in the above theorem, will be deduced as an immediate corollary of lemma 8.1 mentioned above, what does not follow so immediately are retracts of its analytic complements $\Omega_{H}^{A}$. This is detailed in section 9.
Next, it may be noted that an overarching hypothesis that we have imposed on all domains considered thus far, is their boundedness or equivalently (in our context of balanced domains here), hyperbolicity. To deal with the simplest case of unbounded and non-hyperbolic domains then, we first consider a product domain in which one of the factors is still a bounded domain but the other unbounded. We thought it better to separately record the result for the simplest case of this kind, namely of the form ×$\mathbb{C}\times D$ where $D$ is a bounded planar domain; the characterization of its retracts is laid down in the following.
@@
If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of ×$\mathbb{C}\times D$ is either of the form ×$\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
It must be noted that this also addresses the other implicit assumption in most of our previous considerations, namely that of topological triviality. Indeed, the above theorem allows for domains with a variety of fundamental groups; more such variety is facilitated by the next set of results.
@@
While this article is focussed on characterization of retracts rather than retraction mappings, we thought it noteworthy to record in passing, a complete characterization of all retraction maps on the most special domain among the domains in the above theorem namely, ×$\mathbb{C}\times\Delta$.
@@
Every non-trivial retraction map $F$ of ×$\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.19 and more importantly, includes cases with far greater topological complexity.
@@
Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{N}$ itself. While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$ and we shall show by actual examples that it is possible for non-hyperbolic domains to possess hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely, that they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically. It is also easy to see that the one-dimensional retracts of $\mathbb{C}^{N}$ are biholomorphic to $\mathbb{C}$. For more precise results, we consider $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown). Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 10. On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts; more about this in the final section.
Theorem 1.23 (Shpilrain – Yu).
@@
Acknowledgments: We thank Gautam Bharali, J. Jaikrishnan and Kaushal Verma. The second named author acknowledges the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.

### v7

@@ 1. Introduction
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Let $Z$ be any (non-trivial) retract of $D$ through the origin with the property that $\partial D$ is convex near points where $L:=T_{0}Z$ intersects $\partial D$. Then $Z$ can be realized as the graph of a holomorphic map over the linear subspace $D_{L}$ (which itself is a retract of $D$), taking values in $\bigcup_{p\in\partial D_{L}}F_{\mathbb{C}}(p,\partial D)\cap D_{M}$ where $F_{\mathbb{C}}(p,\partial D)$ is the $\mathbb{C}$-linear face of $\overline{D}$ corresponding to the boundary point $p$ and $M$ is a certain linear subspace complementary to $L$ (and as usual $D_{M}:=D\cap M$).
In proving this theorem, we first derive from an idea from [Abate_Isometry], a general result without any convexity hypothesis, thus going beyond the purview of Vesentini’s result. Namely, the following result wherein we assume boundedness. It must be noted that dropping boundedness leads to the failure 888see proposition 10.1 of section 10. of the result whereas replacing boundedness by the apparently more general condition of Kobayashi hyperbolicity is insignificant here, owing to the already mentioned result of Kodama in [Kodama], according to which any hyperbolic balanced domain is bounded.
@@
Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, $Z$ any (non-trivial) retract passing through the origin with $\rho:D\longrightarrow Z$, a retraction map. Let $L$ denote the tangent space to $Z$ at the origin. Then $\rho_{|_{D_{L}}}$ is a biholomorphism mapping $D_{L}=D\cap L$ onto $Z$ and upto a linear change of coordinates, $Z$ is the graph of a holomorphic map over $D_{L}$ taking values in $D_{M}:=\ker\left(D\rho(0)\right)$.
It follows immediately from this theorem that $Z$ is biholomorphic to a (lower dimensional) the linear retract $D_{L}$. In the cases where $\partial D$ is holomorphically extreme, $Z=D_{L}$. Whatever be the case for $D$ and $Z$ as above, several inheritance properties extending the list mentioned earlier follow: if $D$ is $\mathbb{C}$-extremal analytic polyhedron or, strongly pseudoconvex or more generally (pseudoconvex and) of finite type, then $Z$ is biholomorphic to a domain of the same kind. While it is true that when $D$ is an $\ell^{p}$-ball, $Z$ is biholomorphic to a lower dimensional $\ell^{p}$-ball, this is nowhere close to being as trivial as the just mentioned inheritance properties. Next, we note that it does not follow that any two retracts of the same dimension through the origin, are biholomorphically equivalent; indeed, even a pair of linear retracts of the same dimension may fail to be biholomorphic. To mention a concrete example, we may consider $D=\mathbb{B}^{2}\times\Delta^{2}$ wherein both factors can be seen as two-dimensional retracts through the origin in $D$. Relegating further relevant observations here to section 3, we record here a corollary for an important class of domains. Indeed, when $D$ as in the above theorem (without any extremality condition about any boundary point), is homogeneous i.e., ${\rm Aut}(D)$ acts transitively on $D$, then of-course we may glean a characterization of all retracts of such a domain.
@@
Owing to the assumption of being balanced, we recall that the domains as in this corollary are same as the (much studied) bounded balanced symmetric domains which are biholomorphic to (Cartesian) products of the standard Cartan domains of types I – IV, all of whose automorphism groups are well-known (Hua [Hua]); this also means that domains as in the above theorem are convex. This is a good place to draw attention of the reader to the fact that a precise determination of even the linear retracts of a (convex) ball – as in theorem 1.6 for the $l^{p}$-ball for $p<\infty$ and proposition 1.7 for the $p=\infty$ case much earlier above, and few other relevant results as in the already mentioned very recent work [GZ] which deals with among other things, Cartan domains of type IV – is in general quite difficult. Moreover, it must be remembered that an orthogonal projection of such $l^{p}$-balls onto a linear subspace in general is not a retract, except for the case $p=2$. As these are classical facts, the significance of linear retracts from the viewpoint of recent research work may need a remark; let us only mention that it was shown by Mok in his fairly recent work [Mok], that every linear totally geodesic submanifold through the origin of any bounded symmetric domain $\Omega$ with respect to a ${\rm Aut}(\Omega)$-invariant Kähler metric (for instance, the Bergman metric) is realizable as a retract just by orthogonal projection. As this has been dealt with deeply in [Mok], we shall proceed forth next, to look at another direction (but still related to the foregoing corollary in a sense as shall be seen), in relaxing the assumptions in Vesentini’s theorem; wherein which, we retain convexity but drop the extremality condition and ask if finer information can be obtained about the retracts than the general one in the foregoing theorem 1.15. Recall that Vesentini’s result fails if we drop the extremality condition even if we retain convexity, as is manifestly seen in the simplest case of polydiscs. This motivates us to first enquire about the retracts of domains which fail to satisfy extremality conditions at most of its boundary points in the simplest of test cases such as domains decomposable into a product of balls with respect to some norms (called polyballs in [Kuczm_Rtrct_polybll]) satisfying an extremality condition, much as the polydisc. Pursuing this line of enquiry has led us to the result laid down next, which generalizes the result for polydiscs due to Heath and Suffridge, to a much larger class of domains and with a simpler proof following the recent work [BBMV-union]. As indicated above, this theme of retracts of polyballs has indeed been considered earlier for instance in [Kuczm_Rtrct_polybll]; it was shown therein that fixed point sets of holomorphic self-maps of such $N$-fold product of balls are always retracts as well, leaving to be desired however, as to what the retracts of polyballs look like. We shall address this here, in the following next main result of this paper.
@@
  • (A)
    Let $B_{1},B_{2}$ be a pair of bounded balanced domains of holomorphy with holomorphically extreme boundaries, possibly in different complex Euclidean spaces. Then, every non-trivial retract of $D=B_{1}\times B_{2}$ through the origin is the graph, either of a $B_{2}$-valued holomorphic map over a (complex) linear subspace of $B_{1}$ or, of a $B_{1}$-valued holomorphic map over a linear subspace of $B_{2}$.
    More specifically and precisely, let $B_{1}\subset\mathbb{C}^{N_{1}}$ and $B_{2}\subset\mathbb{C}^{N_{2}}$ be as above i.e., not necessarily convex but pseudoconvex bounded balanced domains with holomorphically extreme boundaries. Let $D=B_{1}\times B_{2}$ and suppose without loss of generality that $\min\{N_{1},N_{2}\}=N_{1}$. Let $\mu_{1},\mu_{2}$ denote the Minkowski functionals of $B_{1},B_{2}$ respectively. Let $Z$ be any non-trivial retract of $D$, $L=T_{0}Z$, the tangent space to $Z$ at the origin, $D_{L}=L\cap D$ and $\partial D_{L}=L\cap\partial D$. Then:
@@
      if $\partial D_{L}\subset\partial B_{1}\times\partial B_{2}$, then $Z$ is a complex linear subspace of $D$ which can be realized as the graph of a complex linear map over the linear subspace $\pi_{1}(D_{L})$ of $B_{1}$:
@@
      maps $B_{1}^{L}:=\pi_{1}(D_{L})$ into $B_{2}$,
@@
      if $\partial D_{L}\subset\partial B_{1}\times\overline{B}_{2}$, then $Z$ is the graph of a $B_{2}$-valued holomorphic map over a linear subspace $(B_{1})_{L_{1}}$ of $B_{1}$, given by $(B_{1})_{L_{1}}:=B_{1}\cap L_{1}$, where $L_{1}=\pi_{1}(L)$.
@@
      if $\partial D_{L}\subset\overline{B}_{1}\times\partial B_{2}$, then $Z$ is the graph of a $B_{1}$-valued holomorphic map over a linear subspace $(B_{2})_{L_{2}}$ of $B_{2}$, given by $(B_{2})_{L_{2}}:=B_{2}\cap L_{2}$, where $L_{2}=\pi_{2}(L)$.
@@
  • (B)
    Suppose now that $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ are bounded balanced domains of holomorphy (i.e., with the holomorphic extremal assumption on their boundaries dropped). Then statements (ii) and (iii) of (A) above, hold true. Statements (i) and (iv) hold true if $Z$ is replaced by $D_{L}$ (thus they hold when $Z$ is linear, in which case $Z=D_{L}$; if $Z$ is not linear, then $Z$ is the graph over $D_{L}$).
As the first assertion about retracts through the origin is already contained in theorem 1.15, the real matter of here is the finer information in the statments (i) – (iv) of (A) which we note are not mutually exclusive; it may also be noted that (ii) and (iii) are actually exhaustive. We further note that it may very well happen that neither $\pi_{1}(Z)$ nor $\pi_{2}(Z)$ is a retract of the respective factors $B_{1},B_{2}$ (see the example in remark 3.20). Next, we note that we cannot trivially extend all of part-(A) to extended products $D=B_{1}\times B_{2}\times\ldots\times B_{N}$ in the following manner even in the special case wherein all the $B_{j}$’s to be bounded balanced convex domains each of whose boundaries satisfy the strict convexity condition. Namely: by replacing one of the factors in the above theorem by a product of (N1)(N-1) of these balls i.e., writing for instance $D=(B_{1}\times B_{2}\times\ldots\times B_{N-1})\times B_{N}$, appealing to the above theorem and using an inductive argument; this is due to the fact that $B_{1}\times B_{2}\times\ldots\times B_{N-1}$ fails to satisfy the extremality condition stipulated in part-(A) as soon as $N>2$. However, such an extension is indeed possible and not difficult either; provided only that we repeat the same arguments multiple times by projecting onto the individual factors $B_{j}$ to harness the strict convexity condition satisfied by each of these factor balls of the product $D$. As the complete formulation of the end result of this, turns out to be lengthy, we shall just lay it down in section 3. Here we record an application of part-(B) of the above theorem (in-fact, its essence namely theorem 1.15, suffices) to: the ‘union problem’ for a nice class of polyballs which is interesting owing to it being a case with some degeneracy; this is based on the ideas of Fornæss - Sibony in their article [Fost]. Recall that the union problem comprises the following: let $M$ be a complex manifold which is the union of an increasing sequence of open subsets $M_{j}$ each of which is biholomorphic to a fixed domain $\Omega\subset\mathbb{C}^{N}$; describe $M$ in terms of $\Omega$.
@@
Take $\Omega$ in the union problem to be any homogeneous bounded balanced domain; as is well-known such a domain is a polyball. If $M$ is non-hyperbolic and the corank of its (infinitesimal) Kobayashi metric is one, then $M$ is biholomorphic to $\Omega_{L}\times\mathbb{C}$ where $\Omega_{L}$ is a linear retract of the polyball $\Omega$.
Next, as most balls fail to be homogeneous, the problem of characterizing retracts of bounded balanced pseudoconvex domains not passing through the origin, is in general more difficult and significant. Towards this, we report the following result.
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$, $\rho:D\to Z$ a (non-trivial) retraction map and $p$ the point in $Z$ which is the image of the origin under $\rho$; so, $Z$ is a retract not necessarily passing through the origin. Suppose that: (i) the slice $D_{L}:=D\cap L$ is a bounded convex domain in $L:=T_{p}Z$ with $\mathbb{C}$-extremal boundary and (ii) $D\rho_{|_{0}}$ is a projection which maps $D$ onto $D_{L}$, where $L=T_{p}Z$. Then $Z=\rho(D)$ is the graph of a holomorphic map over the linear subspace $D_{L}$ of $D$.
Examples wherein all the hypotheses of the theorem are satisfied will be provided in the section where we lay down the proof. To move forward, we note that we have for the most part focussed attention on bounded balanced pseudoconvex domains which being contractible, may be regarded as being topologically trivial; moreover, they have other special properties such as Stein neighbourhood bases, and plurisubharmonic barriers all along their boundary (i.e., are hyperconvex) as soon as the Minkowski function is continuous. To turn next to topologically non-trivial cases failing to be hyperconvex, one simple case to begin with in this direction, is that of analytic complements of domains with a Stein neighbourhood basis; specifically, see the remarks following lemma 8.1. Next, to consider the simplest domain lacking a Stein neighbourhood basis and plurisubharmonic barriers throughout the boundary then, we are led to considering the standard domain in this context namely the Hartogs triangle given by
@@
We also contend that this is perhaps the simplest topologically non-trivial case to consider because it is biholomorphic to the product $\Delta\times\left(\Delta\setminus\{0\}\right)$; indeed the map $\phi(z,w)=(w/z,z)$, effects the aforementioned biholomorphic correspondence. Thus it suffices to determine the retracts of $\Delta\times\Delta^{*}$, which we do in the following theorem giving a complete characterization of all retracts of the Hartogs triangle as well as its ‘analytic complements’; this will be deduced as a consequence of a far more general lemma about retracts of ’co-analytic’ domains namely lemma 8.1.
@@
Next, it may be noted that an overarching hypothesis that we have imposed on all domains considered thus far, is their boundedness or, equivalently in our major context of balanced domains here, hyperbolicity. To deal with the simplest case of unbounded and non-hyperbolic domains then, we first consider a product domain in which one of the factors is still a bounded domain but the other unbounded. We thought it better to separately record the result for the simplest case of this kind, namely of the form $\mathbb{C}\times D$ where $D$ is a bounded planar domain; the characterization of its retracts is laid down in the following.
@@
If $D$ is a bounded domain in $\mathbb{C}$, then every non-trivial retract of $\mathbb{C}\times D$ is either of the form $\mathbb{C}\times\{c\}$ for some $c\in D$ or, of the form $\{\left(f(w),w\right)\;:\;w\in D\}$ for some holomorphic function $f$ on $D$; and, conversely every such graph is a retract.
The above theorem allows for domains with a variety of fundamental groups; more such variety is facilitated by the next set of results.
@@
While this article is focussed on characterization of retracts rather than retraction mappings, we thought it noteworthy to record in passing, a complete characterization of the retraction maps on the most special domain among the domains in the above theorem namely, $\mathbb{C}\times\Delta$.
@@
Every non-trivial retraction map $F$ of $\mathbb{C}\times\Delta$, upto conjugation by an automorphism (so that $F$ fixes the origin), is precisely: either of the form $F(z,w)=\left(z+wh(z,w),0\right)$ where $h\in\mathcal{O}(\mathbb{C}\times\Delta)$, or of the form $F(z,w)=(h(w),w)$ where $h\in\mathcal{O}(\Delta)$.
While this article also focusses attention for the most part on domains in $\mathbb{C}^{N}$, we would like to interject here a result about retracts of manifolds, as it is a by-product of the essential ideas in the proof of the foregoing theorem 1.21 and includes cases with far greater topological complexity. For instance, fundamental groups of planar domains are countably generated and in fact always freely generated, whereas for Riemann surfaces this is in general not the case.
@@
It must be noted that this theorem does not subsume proposition 1.22 above, for both factors of the product therein may very well be hyperbolic. We hasten to mention that there are indeed similar ideas in the proofs of the foregoing four results on products, which are therefore dealt with together in section 9 towards reducing repetition of details.
Finally, as one concrete but ‘extreme’ example of an unbounded domain, we may simply take the domain to be $\mathbb{C}^{N}$ itself. While it is rather hard to precisely characterize all retracts of $\mathbb{C}^{N}$, and we shall exemplify the possibility of non-hyperbolic domains possessing hyperbolic retracts, it is easy to observe a qualitative general feature common to all retracts of $\mathbb{C}^{N}$ apart from their topological triviality namely: they are non-hyperbolic to the same (maximal) extent as $\mathbb{C}^{N}$ i.e., their Kobayashi metric vanishes identically. It is also easy to see that the one-dimensional retracts of $\mathbb{C}^{N}$ are biholomorphic to $\mathbb{C}$. For more precise results, we consider $\mathbb{C}^{2}$ – indeed, this is the simplest of balanced domains of holomorphy or for that matter, the simplest of domains with non-trivial retracts – and ask for a complete characterization of all of its holomorphic retracts (which seems unknown). Despite the apparent simplicity of the domain, this seems rather non-trivial; indeed, while graphs (in $\mathbb{C}^{2}$) of entire functions on $\mathbb{C}$ are retracts and the fact that retracts of products of planar domains in the foregoing, have literally turned out to be graphs of holomorphic maps over one of the factors, it is not the case that: every retract of $\mathbb{C}^{2}=\mathbb{C}\times\mathbb{C}$ is the graph of some entire function on either of its factors, as we shall show by providing a concrete example in the last section 10. On the other hand, it is possible to give a complete characterization of polynomial retracts of $\mathbb{C}^{2}$ and this in essence is attained by Shpilrain and Yu in [Shpilrain--Yu] wherein it is deduced from other known results about polynomial rings and their retracts, specifically, the Abhyankar – Moh theorem and a result due to Costa in [Cost]. One may even wonder as to why is this result not a direct corollary of the Abhyankar – Moh theorem; we address this in the final section. We shall however give a self-contained and elementary proof not relying on such theorems, which does not seem to be recorded in the literature to the best of our knowledge; we believe such a proof will help clarify where exactly such theorems are actually necessary. We emphasize for one last time that this result is not new, and is due to various major contributors as clarified above.
Theorem 1.25.
@@
Acknowledgments: We thank Gautam Bharali, C. R. Jayanarayanan, Rohith Varma, J. Jaikrishnan and Kaushal Verma. The second named author would like to acknowledge the support of the CSIR-JRF fellowship provided by the Council of Scientific and Industrial Research, India.